# Projective rigidity of point-line configurations in the plane

### v1

@@
Abstract.
In this paper, we establish a general setup for studying incidence-preserving motions of projective geometric configurations of points and lines via a “projective rigidity matrix”. The spaces of infinitesimal motions of a point-line configuration and dependencies amongst the point-line incidences can be interpreted as the kernel and co-kernel of this projective rigidity matrix, respectively. We also introduce a symmetry-adapted projective rigidity matrix for analysing symmetric configurations and their symmetry-preserving motions. The symmetry may be a point group or a more general symmetry, such as an autopolarity.
Keywords: infinitesimal rigidity; point-line configurations; projective transformations; projective motions; projective self-stresses; symmetry.
1. Introduction
The goal of this paper is to establish new tools, based on methods from geometric rigidity theory [8, 17], for analysing projective configurations of points and lines, that is, collections of points and lines in the real projective plane. This is a classical and still active topic of geometry which dates back to work of T. Kirkman, T. Reye, E. Steinitz, S. Cohn-Vossen and D. Hilbert, among others [14, 15, 10]; see also B. Grünbaum’s book [12]. Projective configurations are also relevant to a variety of problems in practical applications, such as computer vision and architectural design.
A projective configuration defines an incidence relation: a point and a line are incident if the point is on the line. Note that incidences are preserved by projective transformations. In the projective plane every pair of points span a line and every pair of lines meet in a point. Therefore, the question of existence and mobility of projective configurations only becomes interesting when one requires the number of points on each line and the number of lines through each point to exceed $2$. The configuration space of the incidence relation of a projective configuration is the space of all realizations of it in terms of geometric points and lines modulo projective transformations. A given realization is rigid if it is an isolated point of the configuration space, and flexible otherwise. Determining rigidity of a realization is difficult in general. In this article we introduce a new tool which we call the “projective rigidity matrix”. The kernel of this matrix consists of the so-called infinitesimal motions of the configuration, which thereby can be studied using linear algebra. These motions may be thought of as initial velocities of continuous incidence-preserving motions of the points and lines. The co-kernel of the matrix is the space of row-dependencies or self-stresses. It turns out that incidence theorems such as Pappus or Desargues give rise to such row-dependencies/self-stresses; indeed the projective rigidity matrix provides a new tool for studying incidence theorems concerned with points and lines in which one incidence is implied by other incidences.
A symmetry of a projective configuration is a projective transformation that leaves the configuration invariant. Similarly, an auto-polarity is a polarity of the projective plane that leaves the configuration invariant. The study of projective configurations with symmetries and auto-polarities is a classical topic in geometry, see for example [10, 11]. When a configuration moves in the configuration space the symmetries and auto-polarities may or may not be preserved. To study this, we present a symmetry-adapted projective rigidity matrix, called the projective orbit rigidity matrix (based on a similar tool from geometric rigidity theory [16]), whose kernel consists of symmetry-preserving/auto-polarity-preserving infinitesimal motions. To illustrate its usefulness, we apply this new matrix to some examples.
@@ 1. Introduction
The paper is organised as follows. In Section 2 we introduce the necessary background on projective geometry and projective configurations of points and lines. Section 3 establishes the notions of (continuous) rigidity and infinitesimal rigidity for projective configurations. The relationship between these notions is discussed in Section 4. Section 5 is dedicated to the row dependencies of the projective rigidity matrix, which can be interpreted as “self-stresses” of the configuration. Finally, in Section 6 we consider symmetric configurations and set up the projective orbit rigidity matrix for studying symmetry-preserving motions. We conclude the paper with a discussion on future avenues of research in Section 7.
2. Projective configurations
2.1. The projective plane
2.1.1. The analytic model
The real projective plane can be constructed from $\mathbb{R}^{3}$ by letting the $1$-dimensional subspaces of $\mathbb{R}^{3}$ define the points and the $2$-dimensional subspaces define the lines. This construction is called projectivization of the vector space. The homogeneous coordinates of a point in the projective plane are $(x_{0}:y_{0}:z_{0})$ where $(x_{0},y_{0},z_{0})$ is any vector in the corresponding $1$-dimensional subspace. The homogeneous coordinates of a line are $(a_{0}:b_{0}:c_{0})$ where $a_{0}x+b_{0}y+c_{0}z=0$ is any (homogeneous) linear equation defining the corresponding $2$-dimensional subspace. A point with homogeneous coordinates $(x_{0}:y_{0}:z_{0})$ is incident to (or on) a line with homogeneous coordinates $(a_{0}:b_{0}:c_{0})$ if $a_{0}x_{0}+b_{0}y_{0}+c_{0}z_{0}=0$.
The projective plane has a duality; there is a one-to-one correspondence between the points and the lines that preserves the incidences. In terms of vector spaces, a projective duality is a map that sends a $1$-dimensional subspace of $\mathbb{R}^{3}$ to a $1$-dimensional subspace of the dual vector space. A vector $(a_{0},b_{0},c_{0})$ of the dual vector space is a linear form, that is, a linear map from $\mathbb{R}^{3}$ to $\mathbb{R}$ sending $(x,y,z)\mapsto a_{0}x+b_{0}y+c_{0}z$. Each linear form $(a_{0},b_{0},c_{0})$ defines a line of the projective plane, by requiring $a_{0}x+b_{0}y+c_{0}z=0$.
2.1.2. The spherical model
The projective plane can also be defined as the quotient of the unit sphere in $3$-space under identification of antipodal points. In this model, the points of the projective plane are the pairs of antipodal points on the sphere, and the lines in the projective plane are the great circles on the sphere. A great circle on the sphere is the intersection between a plane containing the origin and the sphere.
A duality of the projective plane in this model is defined as a map sending pairs of points to great circles and great circles to pairs of points, preserving incidence. The “standard” duality sends a pair of antipodal points $(p_{+},p_{-})$ to the great circle obtained by intersecting the sphere with the plane through the origin with the position vectors of the pair of points $(p_{+},p_{-})$ as a normal, and it sends a great circle to the pair of antipodal points obtained as the intersection of the sphere with a line through the origin and with direction vector a normal vector of the plane.
2.1.3. The affine plane
Let $l_{\infty}$ be any line of the projective plane and call it the line at infinity. The set of all the points outside $l_{\infty}$ (the finite points) together with the set of all the lines except $l_{\infty}$ defines an affine plane. We select $l_{\infty}$ to be the line defined by the equation $z=0$. Then the homogeneous coordinates of a finite point $(x^{\prime}_{0}:y^{\prime}_{0}:z_{0})$ can be normalized to $(x_{0}:y_{0}:1):=(\frac{x^{\prime}_{0}}{z_{0}}:\frac{y^{\prime}_{0}}{z_{0}}:1)$, where $(x_{0},y_{0})$ are the affine coordinates of the finite point. The projective dual of a finite point $(x_{0}:y_{0}:1)$ is a line with equation $x_{0}x+y_{0}y+1=0$. The affine plane is not preserved by the projective duality; all points have a dual line, but not all lines are duals of points. Specifically, lines through the origin of the affine plane are mapped to points of the line at infinity. By selecting another line to be $l_{\infty}$, one obtains another affine plane, but all these affine planes are isomorphic.
2.2. Plane projective configurations of points and lines
@@
A plane projective configuration is a collection of points and lines in the projective plane together with the incidence relation defined by symmetrized inclusion. In this article all configurations will have a finite number of points and lines.
@@
Any plane projective configuration has an underlying combinatorial object $S=(P,L,I)$ consisting of the set of points $P$, the set of lines $L$ and the incidences $I$ between $P$ and $L$. This combinatorial object is a hypergraph, also known as an incidence geometry of rank 2.
@@
The term combinatorial $(p_{r},l_{k})$-configuration is also used in the literature to denote the underlying combinatorial object of a geometric configuration of $p$ points and $l$ lines for which there are numbers $r$ and $k$ such that $r$ lines go through each point, and $k$ points are on each line. Note that the number of incidences in this case is $|I|=pr=lk$ and the combinatorial configuration is an $r$-regular and $k$-uniform hypergraph with the property that no two hyperedges share more than one vertex. If $r=k$, the configuration is called symmetric (sometimes balanced) and is referred to as a $v_{k}$-configuration if it has $v$ points.
@@
Instead of beginning with the geometric configuration, one can begin with the combinatorial object and study the geometric configurations realizing it. Let $S=(P,L,I)$ be an incidence geometry (a hypergraph) with vertex set $P$ (the elements of which we will call points), hyperedge set $L$ (the elements of which we will call lines) and $I$ a set of point-line incidences. A configuration of points and lines $(S,\mathbf{p},\mathbf{l})$ realizing $S$ is an assignment of a point $\mathbf{p}_{j}=(x_{j}:y_{j}:z_{j})$ in the projective plane to each element $p_{j}$ of $P$, and an assignment of a line $\mathbf{l}_{i}=(a_{i}:b_{i}:c_{i})$ in the projective plane to each element $l_{i}$ of $L$, such that $\mathbf{l}_{i}\cdot\mathbf{p}_{j}=0$ whenever $p_{j}$ and $l_{i}$ are incident. The equation $\mathbf{l}_{i}\cdot\mathbf{p}_{j}=0$ encodes that the point $\mathbf{p}_{j}$ lies on the line $\mathbf{l}_{i}$.
3. Rigidity and flexibility of projective configurations
@@
The configuration space $V(S)$ of an incidence geometry $S=(P,L,I)$ is the space of all collections of assignments $(\mathbf{p},\mathbf{l})$ of points and lines in the projective plane such that $(S,\mathbf{p},\mathbf{l})$ is a configuration of points and lines realizing $S$. In other words, the configuration space is the real projective variety defined by the $|I|$ polynomials of the form $\mathbf{l}_{i}\cdot\mathbf{p}_{j}=0$. Let $V_{\mathbf{p},\mathbf{l}}(S)$ be the subvariety of $V(S)$ consisting of the realizations of $S$ that can be obtained from $(S,\mathbf{p},\mathbf{l})$ by a projective transformation.
@@
Realizing an incidence geometry $S=(P,L,I)$ it can happen that some combinatorial lines or combinatorial points are coincident in the realization, i.e. $\mathbf{l}_{j}=\mathbf{l}_{k}$ for some pair of distinct combinatorial lines $l_{j}$ and $l_{k}$, or $\mathbf{p}_{i}=\mathbf{p}_{m}$ for some pair of distinct combinatorial points $p_{i}$ and $p_{m}$. Note that the configuration space includes all configurations realizing the incidence geometry, including realizations with coincident lines or points.
@@
We say that a configuration $(S,\mathbf{p},\mathbf{l})$ is (projectively) rigid if all configurations $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ in a neighborhood of $(S,\mathbf{p},\mathbf{l})$ in the configuration space are in $V_{\mathbf{p},\mathbf{l}}(S)$. A configuration that is not rigid is said to be (projectively) flexible.
@@
There are several other possible definitions of rigidity for projective configurations of points and lines. One natural definition is as follows: A configuration $(S,\mathbf{p},\mathbf{l})$ is (projectively) rigid if all continuous paths $(\mathbf{p}(t),\mathbf{l}(t))$ in the configuration space such that $\mathbf{p}(0)=\mathbf{p}$ and $\mathbf{l}(0)=\mathbf{l}$ remain in $V_{\mathbf{p},\mathbf{l}}(S)$. With this definition, a configuration $(S,\mathbf{p},\mathbf{l})$ is flexible if there is a continuous path $(\mathbf{p}(t),\mathbf{l}(t))$ in $V(S)$ such that $\mathbf{p}(0)=\mathbf{p}$, $\mathbf{l}(0)=\mathbf{l}$, and such that there is some $t\in(0,1]$ for which $(S,\mathbf{p}(t),\mathbf{l}(t))$ is not in $V_{\mathbf{p},\mathbf{l}}(S)$.
@@
Now, if $(b)$ holds, then there is some largest $t_{0}$ so that $(\mathbf{p}(t_{0}),\mathbf{l}(t_{0}))$ is in $V_{\mathbf{p},\mathbf{l}}(S)$. By definition, there is a projective transformation $A$ that takes $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ to $(S,\mathbf{p},\mathbf{l})$. Every neighbourhood of $(S,\mathbf{p},\mathbf{l})$ intersects the curve $(S,A\mathbf{p}(t),A\mathbf{l}(t))$, for $t\in(t_{0},1]$, and every such intersection necessarily contains configurations not in $V_{\mathbf{p},\mathbf{l}}(S)$, by definition of $t_{0}$. It follows that $(S,\mathbf{p},\mathbf{l})$ is flexible.
It remains to show that $(a)$ implies $(d)$. If $(S,\mathbf{p},\mathbf{l})$ is not rigid, then there is some realization $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ in $U\cap(V(S)\setminus V_{\mathbf{p},\mathbf{l}}(S)$, where $U$ is some neighborhood of $(S,\mathbf{p},\mathbf{l})$ in $V(S)$. By [[18], Lemma 18.3], there are analytic curves $\mathbf{p}(t)$ and $\mathbf{l}(t)$ such that $\mathbf{p}(0)=\mathbf{p}$ and $\mathbf{p}(1)=\mathbf{p^{\prime}}$, and $\mathbf{l}(0)=\mathbf{l}$ and $\mathbf{l}(1)=\mathbf{l^{\prime}}$. Since $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ cannot be obtained from $(S,\mathbf{p},\mathbf{l})$ by a projective transformation, it follows that $(a)$ implies $(d)$. ∎
@@
The motions of interest in this paper are the motions that preserve incidences of points and lines modulo projective transformations. Pinning is a technique that is useful to illustrate the congruence classes of motions. When we pin a configuration we fix four points in general position, that is, in a position such that no three of the points are collinear. Pinning a configuration will leave only the motions that we are interested in.
Pinning a configuration will leave only the motions that we are interested in. We refer the reader to [6] for more details.
The geometric realization of a finite line $l_{i}$ is given by its homogeneous coordinates $\mathbf{l}_{i}=(a_{i}\colon b_{i}\colon 1)$, and the realization of a finite point $p_{j}$ is given by $\mathbf{p}_{j}=(x_{j}\colon y_{j}\colon 1)$. The constraint equation
(1) $a_{i}x_{j}+b_{i}y_{j}+1=0$
@@
Suppose we have a path in the configuration space, and so regard the vectors $\mathbf{p_{j}}$ and $\mathbf{l_{i}}$ as differentiable functions of parameter $t$, say time, whose initial position corresponds to an initial configuration. Since the constraint equations must be satisfied for all values of $t$, the Jacobian of Equation (1) gives a linear equation
(2) $\mathbf{p}_{j}\cdot\Delta\mathbf{l}_{i}+\mathbf{l}_{i}\cdot\Delta\mathbf{p}_{j}=0$
for each incidence $(p_{j},l_{i})\in I$. The coefficient matrix of the system of equations of the form (2) looks as follows:
[…]
A (projective) infinitesimal motion of a configuration $(S,\mathbf{p},\mathbf{l})$ is an element of the kernel of $M(S,\mathbf{p},\mathbf{l})$. Linearizations of projective transformations are infinitesimal motions of all configurations of points and lines. We say that a configuration is (projectively) infinitesimally rigid if all its infinitesimal motions are linearizations of projective transformations. A configuration that is not infinitesimally rigid is (projectively) infinitesimally flexible.

### v3

@@
Abstract
In this paper, we develop tools for analyzing incidence-preserving (or projective) motions of point-line configurations in the real projective plane. We define a real projective variety associated with a rank 3 matroid, that is closely related to the realization space of the matroid. Projective motions are described as curves within this variety, and we introduce the projective rigidity matrix to study these motions. The kernel of this matrix characterizes infinitesimal motions of a configuration, while its co-kernel identifies self-stresses. Whenever part of the matroid corresponds to the points and lines involved in certain projective theorems, such as Desargues’ classical theorem, then the rank of the projective rigidity matrix is reduced, enabling the use of our rigidity matrix for the study of such theorems. Since symmetry often plays a significant role in the construction of point-line configurations, we also introduce a symmetry-adapted projective rigidity matrix for analyzing symmetry-preserving motions, including Euclidean, projective, and autopolarity-preserving motions.
1 Introduction
The study of realization spaces of matroids is a rich and challenging topic, with deep connections to algebraic geometry, combinatorics, and geometric rigidity. For matroids of rank 3, realization spaces capture the possible geometric arrangements of points and lines in the projective plane consistent with the combinatorial structure of the matroid. These spaces have been a focal point in matroid theory, with applications to for example oriented matroids and problems in computational geometry [7, 9]. Realization spaces of matroids are also fundamental in optimization: matroid realization spaces are semialgebraic sets, and Mnëv’s universality theorem states that any semialgebraic set defined over the integers is stably equivalent to the realization space of an oriented matroid [18]. Understanding the realizability of matroids over a given field is a central question, yet it is highly nontrivial. Mnëv’s universality theorem implies that realization spaces of matroids can exhibit arbitrarily complicated topological and geometric behavior. In a recent paper, it was shown that the matroid realization spaces of $\mathbb{C}$-realizable matroids of rank $3$ with a ground set of size 11 or less is smooth, and that for $n\geq 12$, there is a rank $3$ matroid with ground set of size $n$, such that its realization space over $\mathbb{C}$ has singularities [13].
From a geometric perspective, the study of real realizations of rank $3$ matroids can be naturally framed in terms of projective configurations of points and lines in the real projective plane. Projective configurations are classical objects in geometry, with roots in the work of T. Kirkman, T. Reye, E. Steinitz, S. Cohn-Vossen, and D.Hilbert [17, 21, 14]. These configurations define incidence relations: a point is incident with a line if it lies on the line, and every pair of points spans a line while every pair of lines meets at a point. Such structures are preserved under projective transformations. While simple configurations are easily constructed, more intricate configurations, particularly those with higher incidence counts at each object, pose a significant challenge.
Projective configurations offer a unifying framework for bridging geometric and combinatorial approaches to rank $3$ matroids. In matroid terms, the flats correspond to lines, and the bases are triples of points not all collinear. The realization space of such a matroid can be studied through two algebraic frameworks: the matroid realization space and the more recently introduced slack realization space.
A useful computational tool for working with matroid realization spaces is the new computer algebra system OSCAR [19]. See for example [12] for how OSCAR can be applied to matroid realization spaces.
In this paper, we extend the algebraic and geometric study of realization spaces by introducing new tools that enable the analysis of projective configurations of points and lines through the lens of rigidity theory. We define a real projective variety associated with a rank 3 matroid, that contains all points of the classical realization space of the matroid, but also points that correspond to degenerate realizations. Projective motions are described as curves within this variety, and we introduce the projective rigidity matrix to study these motions. The projective rigidity matrix is itself a real realization of the corresponding projective rigidity matroid with ground set the set of incidences between the points and the lines of the configuration. The projective rigidity matrix provides an efficient linear algebraic framework for studying projective configurations. Specifically:
  • Kernel: The kernel of the matrix corresponds to infinitesimal motions of the configuration, which represent initial velocities of continuous, incidence-preserving deformations.
  • Co-kernel: The co-kernel identifies self-stresses, or row dependencies, which are algebraic signatures of implied incidences. For example, classical incidence theorems like Pappus and Desargues manifest as specific self-stresses.
Symmetry, a fundamental aspect of algebraic combinatorics, plays a prominent role in projective configurations. A symmetry of a configuration is a projective transformation that preserves the incidence structure, while an auto-polarity is a projective polarity that leaves the configuration invariant. Motivated by the question whether symmetry can be the underlying reason of existence of certain projective configurations, we introduce the projective orbit rigidity matrix. We prove that the kernel of this matrix consists of the infinitesimal motions that preserve symmetries or auto-polarities, thereby providing a refined tool for studying symmetric configurations.
To illustrate the utility of these matrices, we apply them to various examples, demonstrating their ability to detect rigidity, flexibility, and symmetry-preserving motions in projective configurations. These tools also offer new insights into the realization spaces of matroids, particularly in the presence of symmetry.
@@ 1 Introduction
The paper is organized as follows. Section 2 introduces key concepts in projective geometry and projective configurations. Section 3 develops the notion of (continuous and infinitesimal) rigidity for projective configurations. Section 4 explores self-stresses as row dependencies of the projective rigidity matrix, with connections to classical incidence theorems. Section 5 focuses on symmetric configurations and introduces the projective orbit rigidity matrix for studying symmetry-preserving motions. Finally, Section 6 concludes with a discussion of future directions.
2 Preliminaries
@@
Let $PG(\mathbb{R},2)$ be the real projective plane. Denote by $(x:y:z)$ and $(a:b:c)$ the homogeneous coordinates of a point and a line in $PG(\mathbb{R},2)$, respectively. We call points with $z=0$ points at infinity and lines with $c=0$ are called lines through the origin. A projective configuration is a collection of points and lines in $PG(\mathbb{R},2)$, together with the incidence relation defined by symmetrized inclusion. This definition includes infinite configurations, but in this work we only consider configurations with a finite number of points and lines.
@@
Any projective configuration has an underlying combinatorial object $(P,L,I)$ consisting of the set of points $P$, the set of lines $L$ and the incidence relation $I$ between $P$ and $L$. This is an incidence geometry of rank 2 since it has two types of objects. Such an incidence geometry $(P,L,I)$ can also be regarded as a bipartite graph with vertex set $P\cup L$ and edge set $I$, or as a hypergraph with vertex set $P$ and hyperedge set $L$.
Also, given a rank $2$ incidence geometry, the triples of points that are not incident to a common line are the bases of a rank 3 matroid (the two notions of rank are unrelated).
A flat of a matroid with ground set $E$ is a subset $F\subseteq E$ of the ground set such that the rank of $F\cup\{e\}$ is larger than the rank of $F$ for any element $e\in E\setminus F$. One of many cryptomorphic ways to define a matroid is by its set of flats: a family $\mathcal{F}$ of subsets of $E$ is the set of flats of a matroid if and only if
  1. 1.
    $E\in\mathcal{F}$,
  2. 2.
    If $F_{1}\in\mathcal{F}$ and $F_{2}\in\mathcal{F}$, then $F_{1}\cap F_{2}\in\mathcal{F}$, and
  3. 3.
    If $F\in\mathcal{F}$ and $\{F_{1},F_{2},...,F_{k}\}$ is a minimal set of members of $\mathcal{F}$ that properly contain $F$, then the sets $F_{1}\setminus F$, $F_{2}\setminus F$,…, $F_{k}\setminus F$ partition $E\setminus F$.
The lines of an incidence geometry, and the pairs of points that are not incident to a common line of the incidence gemetry, are the flats of rank 2 of the rank 3 matroid defined by the incidence geometry. See for example one of the textbooks by Oxley or Björner et al. for more about matroids and matroid realizations [7, 20].
From a geometric perspective, it is interesting to study the realization spaces of rank $3$ matroids. There are two algebraic models of the realization space.
First, we describe the matroid realization space. Let $M$ be a rank $3$ matroid. Assign the variables $p_{i}=(x_{i},y_{i},z_{i})$ to the element $p_{i}$ of the ground set of $M$, and let $[p_{i}p_{k}p_{j}]$ denote the determinant of the $3\times 3$-matrix with column vectors $p_{i}$, $p_{j}$ and $p_{k}$. The matroid realization space of $M$ is the subset of the Grassmannian of the lines of the projective plane defined by the following:
$\begin{cases}[p_{i}p_{k}p_{j}]=0&\quad\text{if }\{p_{i},p_{j},p_{k}\}\text{ isnot a basis of }M\\ [p_{i}p_{k}p_{j}]\neq 0&\quad\text{if }\{p_{i},p_{j},p_{k}\}\text{ is a basis of }M.\end{cases}$
This definition can be found for example in work by Björner et al. [7].
In contrast, the slack realization space, recently introduced by Brandt and Wiebe [10], is defined from the symbolic slack matrix $S_{M}(x)$; for a rank $3$ matroid $M$ this is a matrix with rows indexed by the points, and columns indexed by the lines. The entry $S_{M}(x)_{ij}$ is 00 if the element $p_{i}$ lies on the line $l_{j}$, and a variable $x_{ij}$ if the point $p_{i}$ does not lie on the line $l_{j}$. Let $t$ be the number of variables in $S_{M}(x)$, and let $J$ be the ideal in $\mathbb{R}[x_{1},...,x_{t}]$ defined by the $4\times 4$-minors of the slack matrix. The slack ideal $I_{M}$ is the saturation of $J$ with respect to the principal ideal generated by the product of all variables $x_{ij}$. The slack realization space is the variety $V(I_{M})$ in $\mathbb{R}^{t}$ defined by the slack ideal. The matrix $S_{M}(s)$ realizes the matroid $M$ if and only if $s\in I_{M}$ [10].
@@
In the literature, the term combinatorial $(p_{r},l_{k})$-configuration is used to denote a purely combinatorial incidence geometry of rank 2 with the following properties: (i) there are $p$ “points” and $l$ “lines”, (ii) there are positive integers $r$ and $k$ such that $r$ “lines” go through each “point”, and $k$ “points” are on each “line”, and (iii) each pair of “lines” meet in at most one “point” and each pair of “points” span at most one “line”. Note that the number of incidences in this case is $|I|=pr=lk$. In the language of hypergraphs, a combinatorial $(p_{r},l_{k})$-configuration is an $r$-regular and $k$-uniform hypergraph with the property that no two hyperedges share more than one vertex. If $r=k$, then the configuration is called balanced (sometimes symmetric) and is referred to as a combinatorial $v_{k}$ configuration if it has $v$ points.
The literature of projective configurations is often concerned with questions of existence and constructions. Some combinatorial incidence geometries cannot be realized as points and straight lines in the projective plane in such a way that all points and lines are distinct. Given an incidence geometry, one can therefore ask whether it is realizable. In this paper, we start with a realization of a combinatorial incidence geometry, and investigate the projective rigidity properties of that realization.
@@
A (projective planar) realization of an incidence geometry $S=(P,L,I)$ is an assignment of a point $\mathbf{p}_{j}=(x_{j}:y_{j}:z_{j})$ in the projective plane to each element $p_{j}$ of $P$, and an assignment of a line $\mathbf{l}_{i}=(a_{i}:b_{i}:c_{i})$ in the projective plane to each element $l_{i}$ of $L$, such that $\mathbf{l}_{i}\cdot\mathbf{p}_{j}=0$ whenever $p_{j}$ and $l_{i}$ are incident. The equation $\mathbf{l}_{i}\cdot\mathbf{p}_{j}=0$ then encodes that the point $\mathbf{p}_{j}$ lies on the line $\mathbf{l}_{i}$. Note that this definition does not exclude that points or lines coincide.
In Section 5, we will be particularly interested in symmetric realizations, which are realizations that are invariant under the action of a symmetry group. The symmetries that we are considering can be either Euclidean symmetries, for example reflections or rotations, or they can be autopolarities.
3 Rigidity and flexibility of projective configurations
@@
The realization space $V(S)$ of an incidence geometry $S=(P,L,I)$ is the space of all collections of assignments $(\mathbf{p},\mathbf{l})$ of points and lines in the projective plane such that $(S,\mathbf{p},\mathbf{l})$ is a configuration of points and lines realizing $S$. In other words, the realization space is the real projective variety defined by the $|I|$ polynomials of the form $\mathbf{l}_{i}\cdot\mathbf{p}_{j}=0$. Let $V_{\mathbf{p},\mathbf{l}}(S)$ be the subvariety of $V(S)$ consisting of the realizations of $S$ that can be obtained from $(S,\mathbf{p},\mathbf{l})$ by a projective transformation.
Suppose that a rank 3 matroid has as its bases the triples of three points that are not incident to a common line in an incidence geometry. Then the points of the Grassmannian of that matroid, and the points of its slack realization space can be seen as realizations of the incidence geometry as points and lines in $PG(\mathbb{R},2)^{2}$. However, the realizations in the Grassmannian of that rank $3$ matroid and in the slack realization space are exactly those where a realized point lies on a realized line if, and only if, the combinatorial point is incident to the combinatorial line.
@@
The realization space defined above contains all collections of points and lines that satisfy the combinatorial incidences. Therefore, it can happen that some combinatorial lines or combinatorial points are coincident in the realization, i.e. $\mathbf{l}_{j}=\mathbf{l}_{k}$ for some pair of distinct combinatorial lines $l_{j}$ and $l_{k}$, or $\mathbf{p}_{i}=\mathbf{p}_{m}$ for some pair of distinct combinatorial points $p_{i}$ and $p_{m}$.
@@
We say that a configuration $(S,\mathbf{p},\mathbf{l})$ is (projectively) rigid if all configurations $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ in a neighborhood of $(S,\mathbf{p},\mathbf{l})$ in the realization space are in $V_{\mathbf{p},\mathbf{l}}(S)$. A configuration that is not rigid is said to be (projectively) flexible.
@@
There are several other possible definitions of rigidity for projective configurations of points and lines. One natural definition is as follows: A configuration $(S,\mathbf{p},\mathbf{l})$ is (projectively) rigid if all continuous paths $(\mathbf{p}(t),\mathbf{l}(t))$ in the realization space such that $\mathbf{p}(0)=\mathbf{p}$ and $\mathbf{l}(0)=\mathbf{l}$ remain in $V_{\mathbf{p},\mathbf{l}}(S)$. With this definition, a configuration $(S,\mathbf{p},\mathbf{l})$ is flexible if there is a continuous path $(\mathbf{p}(t),\mathbf{l}(t))$ in $V(S)$ such that $\mathbf{p}(0)=\mathbf{p}$, $\mathbf{l}(0)=\mathbf{l}$, and such that there is some $t\in(0,1]$ for which $(S,\mathbf{p}(t),\mathbf{l}(t))$ is not in $V_{\mathbf{p},\mathbf{l}}(S)$.
@@
Now, if $(b)$ holds, then there is some largest $t_{0}$ so that $(\mathbf{p}(t_{0}),\mathbf{l}(t_{0}))$ is in $V_{\mathbf{p},\mathbf{l}}(S)$. By definition, there is a projective transformation $A$ that takes $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ to $(S,\mathbf{p},\mathbf{l})$. Every neighborhood of $(S,\mathbf{p},\mathbf{l})$ intersects the curve $(S,A\mathbf{p}(t),A\mathbf{l}(t))$, for $t\in(t_{0},1]$, and every such intersection necessarily contains configurations not in $V_{\mathbf{p},\mathbf{l}}(S)$, by definition of $t_{0}$. It follows that $(S,\mathbf{p},\mathbf{l})$ is flexible.
It remains to show that $(a)$ implies $(d)$. If $(S,\mathbf{p},\mathbf{l})$ is not rigid, then there is some realization $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ in $U\cap(V(S)\setminus V_{\mathbf{p},\mathbf{l}}(S)$, where $U$ is some neighborhood of $(S,\mathbf{p},\mathbf{l})$ in $V(S)$. Without loss of generality, we can assume that both realizations have no points at infinity or lines that go through the origin.
The realizations that have only finite points, and only lines that do not go through the origin correspond to points on the real affine variety defined by a polynomial of the form
$a_{i}x_{j}+b_{i}y_{j}+1=0$ (1)
for each incidence $(p_{j},l_{i})$.
By [[23], Lemma 18.3], there are analytic curves $\mathbf{p}(t)$ and $\mathbf{l}(t)$ in this real affine variety such that $\mathbf{p}(0)=\mathbf{p}$ and $\mathbf{p}(1)=\mathbf{p^{\prime}}$, and $\mathbf{l}(0)=\mathbf{l}$ and $\mathbf{l}(1)=\mathbf{l^{\prime}}$. The curves $\mathbf{p}(t)$ and $\mathbf{l}(t)$ lift to analyic curves in $V(S)$ such that $z(t)$ and $c(t)$ are constant for all points and lines respectively. Since $(S,\mathbf{p^{\prime}},\mathbf{l^{\prime}})$ cannot be obtained from $(S,\mathbf{p},\mathbf{l})$ by a projective transformation, it follows that $(a)$ implies $(d)$. ∎
@@
The motions of interest in this paper are the motions that preserve incidences of points and lines modulo projective transformations. Pinning is a technique that is useful to illustrate the congruence classes of motions. When we pin a configuration we fix four points in general position, that is, in a position such that no three of the points are collinear. Pinning a configuration will leave only the motions that we are interested in. We refer the reader to [6] for more details.
The geometric realization of a finite line $l_{i}$ that does not go through the origin is given by its homogeneous coordinates $\mathbf{l}_{i}=(a_{i}\colon b_{i}\colon 1)$, and the realization of a finite point $p_{j}$ is given by $\mathbf{p}_{j}=(x_{j}\colon y_{j}\colon 1)$. The constraint equation
$a_{i}x_{j}+b_{i}y_{j}+1=0$ (2)
@@
Suppose we have a path in the realization space, and so regard the vectors $\mathbf{p_{j}}$ and $\mathbf{l_{i}}$ as differentiable functions of parameter $t$, say time, whose initial position corresponds to an initial configuration. Since the constraint equations must be satisfied for all values of $t$, the Jacobian of Equation (2) gives a linear equation
$\mathbf{p}_{j}\cdot\Delta\mathbf{l}_{i}+\mathbf{l}_{i}\cdot\Delta\mathbf{p}_{j}=0$ (3)
for each incidence $(p_{j},l_{i})\in I$. The projective rigidity matrix is the coefficient matrix $M(S,\mathbf{p},\mathbf{l})$ of the system of equations of the form (3). The row of the projective rigidity matrix that corresponds to the incidence $(p_{j},l_{i})$ looks as follows:
$\left[\begin{array}[]{ c ccccc ccccc }&0&\dots&x_{j}\ y_{j}&\dots&0&\ldots&a_{i}\ b_{i}&\ldots&0&\end{array}\right],$
A (projective) infinitesimal motion of a configuration $(S,\mathbf{p},\mathbf{l})$ is an element of the kernel of $M(S,\mathbf{p},\mathbf{l})$. Linearizations of projective transformations are infinitesimal motions of all configurations of points and lines. We say that a configuration is (projectively) infinitesimally rigid if all its infinitesimal motions are linearizations of projective transformations. A configuration that is not infinitesimally rigid is (projectively) infinitesimally flexible. The following relation between rigidity and infinitesimal rigidity holds:
Theorem 2 ([6]).
If a configuration is projectively infinitesimally rigid, then it is projectively rigid.